score is 096. You have unlimited attempts remaining.
Answer #1

Mortgage amount = Cost of the house - Down payment =180000*(1-.20) = 144000

Monthly payments is calculated using the PMT function as follows:

=PMT(rate,nper,pv)

=PMT(5.5%/12,30*12,144000)

=817.62